OpenStudy (anonymous):

It says to graph each linear equation how would i do this: x-2y=6

OpenStudy (anonymous):

write it in the form y = mx + b first 2y = x - 6 y = x/(2) - 3

OpenStudy (anonymous):

the y-intercept is -3, so the point (0, -3) is on the graph. Plot that point, then find one other point on the graph and plot it

OpenStudy (anonymous):

for example, when x = 2, y = -2, so (2, -2) is another point on the line.

OpenStudy (anonymous):

one you've plotted two points, draw a line through them and you're done

OpenStudy (anonymous):

Wait it says you got it wrong

OpenStudy (anonymous):

in the back of my book its plotted as 3-1/2

OpenStudy (anonymous):

err, what does the back say?

OpenStudy (anonymous):

because there are other methods, but this one wasn't wrong.

OpenStudy (anonymous):

Ya, you got it rite, Thank you for all the help man, I really needed it.. Thanks again!

OpenStudy (anonymous):

But how would i solve for this 6x+3=y